### v1

# Controlled LLM Training on Spectral Sphere

Abstract
Scaling large models requires optimization strategies that ensure rapid convergence grounded in stability. Maximal Update Parametrization ($\bm{\mu}$P) provides a theoretical safeguard for width-invariant $\Theta(1)$ activation control, whereas emerging optimizers like Muon are only “half-aligned” with these constraints: they control updates but allow weights to drift. To address this limitation, we introduce the Spectral Sphere Optimizer (SSO), which enforces strict module-wise spectral constraints on both weights and their updates. By deriving the steepest descent direction on the spectral sphere, SSO realizes a fully $\bm{\mu}$P-aligned optimization process. To enable large‑scale training, we implement SSO as an efficient parallel algorithm within Megatron. Through extensive pretraining on diverse architectures, including Dense 1.7B, MoE 8B-A1B, and 200-layer DeepNet models, SSO consistently outperforms AdamW and Muon. Furthermore, we observe significant practical stability benefits, including improved MoE router load balancing, suppressed outliers, and strictly bounded activations.
[Uncaptioned image]Project Page: Spectral-Sphere-Optimizer
((a)) AbsMax (outliers) of Attention Activations
((b)) RMS of FFN Activations
Figure 1: Training dynamics of Dense-1.7B activations (log-scaled cross-layer averages). Our Spectral Sphere maintains constant activation magnitudes throughout training because its $\bm{\mu}$P-metrized constraints on the spectral manifold ensure that the activation RMS remains at $\Theta(1)$ scale. Muon activations show a mild drift due to learning rate decay and weight decay. By contrast, AdamW proves the most unstable, generating significantly larger activations, with attention AbsMax and FFN RMS reaching $\sim\!100\times$ magnitude compared to those spectral optimizers.
Contents
1 Introduction
LLM training is, at its core, a pursuit of convergence speed grounded in the necessity of stability. While the community has explored various optimization strategies, we argue that the essential principle governing training stability is the Maximal Update Parametrization ($\bm{\mu}$P) (yang2023spectral). By mandating that the spectral norms of weights and updates scale as $\Theta(\sqrt{d_{\text{out}}/d_{\text{in}}})$ to ensure width-invariant activations remains $\Theta(1)$ scale, $\bm{\mu}$P serves as the mathematical safeguard against activation explosions (takase2025spikemorestabilizingpretraining). However, the current landscape is saturated with methods that fail to satisfy these fundamental conditions. Conventional soft regularization methods, such as decoupled weight decay or initialization strategies, prove insufficient over long training horizons (kosson2025weightdecaymattermup). This unconstrained weight drift destabilizes the effective step size (update-to-weight ratio) and degrades feature learning.
On the other side of the spectrum lies the pursuit of optimal convergence. The recent Muon optimizer (jordan2024muon) has demonstrated remarkable efficiency, often interpreted as steepest descent under the spectral norm. In analyzing Muon, we uncover a surprising insight: it acts as a “half-aligned” solution to the $\bm{\mu}$P constraints. However, maintaining stable features requires constraining not only the updates but also the weights themselves. Unstable activations like attention logits explosion were still observed in Muon training (kimiteam2025kimik2openagentic). Consequently, practitioners are forced to rely on “non-essential” architectural patches to artificially force stability — ranging from aggressive normalization schemes like SandwichNorm (ding2021cogviewmasteringtexttoimagegeneration) and QK-Norm (henry2020querykeynormalizationtransformers), to ad-hoc fixes like logit softcapping (kimiteam2025kimik2openagentic) — often at the cost of model expressivity and requiring extensive hyperparameter tuning. This observation motivates a fundamental question:
What if an optimizer could simultaneously satisfy the steepest descent property for convergence speed and the strict $\bm{\mu}$P constraints for fundamental stability?
To answer this, we propose the mathematically unique solution that unifies these two objectives. By identifying the spectral sphere as the natural manifold for stable feature learning, SSO derives the steepest descent direction constrained within this geometry. Unlike heuristic manifold projection methods (xie2025mhcmanifoldconstrainedhyperconnections; pethick2025trainingdeeplearningmodels), SSO solves a constrained optimization problem in the tangent space via a Lagrange multiplier search, followed by a retraction step to map the trajectory back onto the spectral sphere.
To enable large-scale training, we offer a systematic implementation in Megatron. We provide principled guidelines for spectral preconditioned optimization, specifically deriving the optimal learning rate scaler, determining the critical atomic module granularity, and identifying the optimal spectral radius to control activation at optimal scales precisely. These offer a robust recipe for large-scale training. Specifically to mitigate the overhead of the iterative root solver, we utilize a novel distributed strategy centered on atomic module sharding(emergeing_optimizer). This technique partitions fused params into independent spectral units, enabling communication-free local updates. We further address solver-induced workload imbalance through a size-aware ping-pong placement strategy, and accelerate matrix operations using adaptive kernel dispatcher, alongside multi-stream execution and singular vector caching.
Empirically, we validate SSO through extensive pretraining experiments across various scales, including Dense 1.7B, MoE 8B-A1B, and 200-layer DeepNet models. SSO consistently outperforms AdamW and Muon while uniquely preserving stable $\bm{\mu}$P learning rate transfer. Notably, it yields superior training dynamics: it significantly improves MoE router load balancing, suppresses outliers in deep networks, and strictly bounds activations within a tunable scale.
2 Preliminary
2.1 Maximal Update Parametrization ($\bm{\mu}$P)
$\bm{\mu}$P prescribes how activations and weight updates should scale with width to preserve feature learning (yang2023spectral). Ideal feature learning requires the scale of activations to remain as invariant as possible. We use operator norm to characterize how the norm of activations changes through a linear layer.
Considering a linear layer ${\bm{y}}={\bm{W}}{\bm{x}}$ with ${\bm{W}}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}},{\bm{x}}\in\mathbb{R}^{d_{\mathrm{in}}}$, the RMS norm is defined as
$\|{\bm{x}}\|_{\mathrm{rms}}:=\frac{\|{\bm{x}}\|_{2}}{\sqrt{d_{\mathrm{in}}}},$ (1)
while the RMS-to-RMS operator norm is defined as
$\|{\bm{W}}\|_{\mathrm{rms}\to\mathrm{rms}}:=\sup_{{\bm{x}}\neq\bm{0}}\frac{\|{\bm{W}}{\bm{x}}\|_{\mathrm{rms}}}{\|{\bm{x}}\|_{\mathrm{rms}}}.$ (2)
$\bm{\mu}$P scale invariance requires maintaining $\|{\bm{y}}\|_{\mathrm{rms}}=\|{\bm{x}}\|_{\mathrm{rms}}=\Theta(1),$ which is equivalent to enforcing the RMS-to-RMS condition $\|{\bm{W}}\|_{\mathrm{rms}\to\mathrm{rms}}=\|{\bm{W}}\|_{2}\sqrt{{d_{\mathrm{in}}}/{d_{\mathrm{out}}}}=\Theta(1).$ This induces the following spectral norm constraint on the weight matrix $\|{\bm{W}}\|_{2}=\Theta(\sqrt{{d_{\mathrm{out}}}/{d_{\mathrm{in}}}}).$ A similar requirement applies to parameter updates; together, we refer to these as the spectral $\bm{\mu}$P condition below.
yang2023spectral shows that preserving scale-invariant activations for feature learning requires the same law to hold for both weights and their updates: $\|{\bm{W}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{int}}}}\right),\qquad\|{\bm{\Phi}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{int}}}}\right).$
2.2 Steepest Descent under Different Norms
Following metrized deep learning (bernstein2024oldoptimizernewnorm), many successful optimizers (e.g. AdamW (loshchilov2019decoupled), Shampoo (gupta2018shampoo), Prodigy (mishchenko2024prodigy)) can be interpreted as first-order methods without convexity assumptions. Specifically, after switching off exponential moving averages, these algorithms reduce to instances of steepest descent governed by distinct norms. Under this framework, an optimizer is fundamentally defined by a weight update $\bm{\Phi}$} that minimizes a quadratic model of the loss:
${\bm{\Phi}}:=\underset{{\bm{\Phi}}}{\operatorname*{argmin}}\left\{\mathcal{L}({\bm{W}})+\langle{\bm{G}},{\bm{\Phi}}\rangle+\frac{\lambda}{2}\|{\bm{\Phi}}\|^{2}\right\}.$ (3)
The update is thus determined by two priors: a norm $\|\!\cdot\!\|$ assigned according to the specific functional role of the module, and a sharpness $\lambda$ that governs the update scale. The solution is as follows:
Given a norm $\|\!\cdot\!\|$ that endows the parameter space with a geometry, the steepest descent update is ${\bm{\Phi}}=-\eta\cdot{\bm{\Phi}}_{\text{unit}},\quad\text{where}\quad{\bm{\Phi}}_{\text{unit}}:=\underset{\|{\bm{T}}\|\leq 1}{\arg\max}\langle{\bm{G}},{\bm{T}}\rangle,\quad\text{and}\quad\eta:=\frac{\|{\bm{G}}\|_{\dagger}}{\lambda}.$ (4) Here, $\|{\bm{G}}\|_{\dagger}:=\max_{\|{\bm{T}}\|\leq 1}\langle{\bm{G}},{\bm{T}}\rangle$ is the dual norm of the gradient $\bm{G}$} induced by $\|\!\cdot\!\|$.
This perspective unifies diverse algorithms by mapping them to their underlying geometries: SGD corresponds to Frobenius norm $\|\!\cdot\!\|_{F}$, Adam to $l_{\infty}$ norm, and Shampoo to spectral norm $\|\!\cdot\!\|_{2}$.
2.3 Muon Optimizer
Following the framework of metrized deep learning (Section˜2.2), Muon (jordan2024muon) can be interpreted as steepest descent under the spectral norm. For a 2D weight matrix $\bm{W}$}, the spectral norm $\|\cdot\|_{2}$  [^1] For vectors, $\|{\bm{v}}\|_{2}$ denotes the $\ell_{2}$ norm; for matrices, $\|{\bm{W}}\|_{2}$ denotes the (induced $\ell_{2}\!\to\!\ell_{2}$) spectral norm. gives the tightest bound on the matrix’s input-output gain:
$\|{\bm{W}}\|_{2}:=\sup_{{\bm{x}}\neq\bm{0}}\frac{\|{\bm{W}}{\bm{x}}\|_{2}}{\|{\bm{x}}\|_{2}},$ (5)
By choosing the spectral norm to constrain the update direction $\bm{\Phi}$}, the steepest descent direction is uniquely given by the matrix sign function (msign) (Section˜A.1):
$\operatorname{msign}({\bm{G}})={\bm{U}}\operatorname{sign}(\bm{\Sigma}){\bm{V}}^{\top}={\bm{U}}_{[:,:r]}{\bm{V}}_{[:,:r]}^{\top},$ (6)
where ${\bm{G}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}$ is the singular value decomposition (SVD). The msign operation orthogonalizes the gradient, equalizing all active singular directions and yielding a spectrum isotropic update. A key contribution of Muon is the efficient approximation of $\operatorname{msign}({\bm{G}})$ via Newton–Schulz iterations which can be implemented on GPUs.
However, Muon constrains only the backward update $\bm{\Phi}$}, leaving the forward weights $\bm{W}$} unconstrained. This often leads to unstable $\bm{\mu}$P feature learning in hidden states rms, motivating our development of a fully aligned approach that constrains both $\bm{W}$} and $\bm{\Phi}$}.
[…]
3 Method
3.1 Optimization Target Formulation
We start by focusing on a hidden layer matrix ${\bm{W}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$. To satisfy the spectral $\bm{\mu}$P scaling condition in Section˜2.1, we set the spectral scale to target radius $R$:
$R=\Theta\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right).$ (7)
Following metrized deep learning (Section˜2.2), we define the unit update $\bm{\Phi}$} as the solution to a constrained optimization problem:
We perform steepest descent under the spectral norm, constraining both the weights and the updates to a spectral sphere of radius $R$. Specifically, we parameterize the update step as $\Delta{\bm{W}}=\eta R{\bm{\Phi}}$, where $\eta$ is the base learning rate. The update direction $\bm{\Phi}$} is the solution to: $\max_{{\bm{\Phi}}}$ $\langle{\bm{G}},{\bm{\Phi}}\rangle$ (8) s.t. $\|{\bm{\Phi}}\|_{2}=1,$ $\|{\bm{W}}-\eta R{\bm{\Phi}}\|_{2}=\|{\bm{W}}\|_{2}=R.$
3.2 First-Order Tangent Space Constraint
Assisted by the uniqueness of the top singular value [^2] Numerical coincidence of singular values is a measure-zero event for random matrices (tao2014randommatricessimplespectrum). Quantitatively, it occurs with probability at most $\exp(-c\max(d_{\mathrm{in}},d_{\mathrm{out}}))$ (han2025repeatedsingularvaluesrandom)., the spectral norm $\|{\bm{W}}\|_{2}$ is differentiable with gradient ${\bm{\Theta}}:=\nabla_{{\bm{W}}}\|{\bm{W}}\|_{2}={\bm{u}}_{1}{\bm{v}}_{1}^{\top}$, where $({\bm{u}}_{1},{\bm{v}}_{1})$ are the principal left and right singular vectors (watson1992characterization). We consider a first-order Taylor Expansion of the spectral norm around $\bm{W}$}:
$\|{\bm{W}}-\eta R{\bm{\Phi}}\|_{2}=\|{\bm{W}}\|_{2}-\eta R\langle{\bm{\Theta}},{\bm{\Phi}}\rangle+\mathcal{O}(\eta^{2}R^{2}\|{\bm{\Phi}}\|_{2}^{2}).$ (9)
[…]
$\max_{{\bm{\Phi}}}\ \langle{\bm{G}},{\bm{\Phi}}\rangle\quad\text{s.t.}\quad\|{\bm{\Phi}}\|_{2}\ =1,\ \ \langle{\bm{\Theta}},{\bm{\Phi}}\rangle=0.$ (10)
We then introduce a Lagrange multiplier $\lambda$ and maximize Lagrangian $\mathcal{L}({\bm{\Phi}},\lambda)=\langle{\bm{G}}+\lambda{\bm{\Theta}},{\bm{\Phi}}\rangle$ under constraint $\|{\bm{\Phi}}\|_{2}=1$. The analytical solution and numerical method are summarized below.
[…]
((a)) $1024\times 3072$ matrix
((b)) $4096\times 1024$ matrix
Figure 3: Empirical curves of $h(\lambda)=\langle{\bm{\Theta}},\operatorname{msign}({\bm{G}}+\lambda{\bm{\Theta}})\rangle$ for random matrices. $h(\lambda)$ is monotonic non-decreasing in $\lambda$, and its root $\lambda^{\star}$ lies close to zero (proof in Section˜A.2). Here, each curve is obtained by averaging over 5 repeats, and each matrix is initialized from $\mathcal{N}(0,0.02^{2})$.
3.3 Second-Order Manifold Constraint
Note that the remainder $\mathcal{O}(\eta^{2}R^{2}\|{\bm{\Phi}}\|_{2}^{2})$ in Equation˜9 may accumulate over iterations, causing gradual drift off the spectral sphere. To enforce the exact constraint $\|{\bm{W}}\|_{2}=R$ throughout training, we apply a retraction step that projects the weights back onto the manifold:
${\bm{W}}\leftarrow{\bm{W}}\cdot\frac{R}{\|{\bm{W}}\|_{2}}.$ (13)
While the retraction is conceptually a post-update projection, we implement it as a pre-update correction for efficiency, which is operationally equivalent. This reordering allows us to invoke the computationally expensive Power Iteration only once per step, reusing the resulting singular triplet for both the manifold retraction and the tangent projector $\bm{\Theta}$} (Lines 6–9 in Algorithm˜1).
The retraction strictly constrains $\|{\bm{W}}\|_{2}=R$, which via Equation˜14, automatically bounds the weight magnitudes. As a result, weight decay, which is primarily introduced to limit the weight scale, becomes redundant. We therefore eliminate weight decay in hidden 2D weights [^4] We still apply weight decay to 1D params (e.g. Embedding, RMSNorm) for possible scaling stability, although our ablations on 1.7B model suggest that disabling 1D params may actually be slightly better., removing a sensitive hyperparameter from training. More details can be found in Section˜A.3.
$\|{\bm{W}}\|_{F}\leq\sqrt{rank({\bm{W}})}\,\|{\bm{W}}\|_{2}\leq\sqrt{\min(d_{\text{out}},d_{\text{in}})}\,R.$ (14)
3.4 Overall Algorithm & Interpretation
Algorithm 1 Spectral Sphere Optimizer (SSO)
1:Initial 2D weights ${\bm{W}}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, spectral $\bm{\mu}$P scaler $R=\sqrt{d_{\text{out}}/d_{\text{in}}}$, learning rate $\eta$, momentum coefficient $\beta$, precision tolerance $\epsilon$
2:Initialize: ${\bm{W}}_{0}\leftarrow R\cdot{\bm{W}}_{0}/\|{\bm{W}}_{0}\|_{2}$, ${\bm{M}}_{0}\leftarrow 0$
3:for $t=0,1,\dots$ do
4:  ${\bm{G}}_{t}\leftarrow\nabla_{\bm{W}}\mathcal{L}({\bm{W}}_{t})$
5:  ${\bm{M}}_{t}\leftarrow\beta{\bm{M}}_{t}+(1-\beta){\bm{G}}_{t}$
6:  $\widehat{{\bm{M}}}_{t}\leftarrow{\bm{M}}_{t}/\|{\bm{M}}_{t}\|_{F}$ $\triangleright$ Normalize for Numerical Stability
7:// 1. Spectral Geometry Analysis
8:  $({\sigma}_{t},{\bm{u}}_{t},{\bm{v}}_{t})\leftarrow\mathrm{PowerIteration}({\bm{W}}_{t})$ $\triangleright$ Top Singular Value & Vectors
9:  ${\bm{\Theta}}_{t}\leftarrow{\bm{u}}_{t}{\bm{v}}_{t}^{\top}$ $\triangleright$ Tangent Space Projector
10:// 2. Retraction to Spectral Sphere
11:  ${\bm{W}}_{t}\leftarrow{\bm{W}}_{t}\cdot R/{\sigma}_{t}$
12:// 3. Steepest Descent Lagrange Solver
13:  Define $h(\lambda)\coloneqq\langle{\bm{\Theta}}_{t},\operatorname{msign}(\widehat{{\bm{M}}}_{t}+\lambda{\bm{\Theta}}_{t})\rangle$
14:  $\lambda_{t}^{*}\leftarrow\mathrm{Bisection}(h,\text{tolerance}=\epsilon)$ $\triangleright$ Find root of $h(\lambda)=0$
15:// 4. $\bm{\mu}$P-Scaled Update
16:  ${\bm{\Phi}}_{t}\leftarrow\operatorname{msign}(\widehat{{\bm{M}}}_{t}+\lambda_{t}^{*}{\bm{\Theta}}_{t})$
17:  ${\bm{W}}_{t+1}\leftarrow{\bm{W}}_{t}-\eta\cdot R\cdot{\bm{\Phi}}_{t}$ $\triangleright$ $\bm{\mu}$P Style Update
18:end for
[…]
Figure 4: Geometry of Steepest Descent Update Directions. The left solid arc denotes the $\bm{W}$} sphere, while the right dotted arc denotes the $\Delta{\bm{W}}$sphere (unit $\bm{\Phi}$} scaled by $\eta$). The shaded region represents the feasible set within the tangent space of the $\bm{W}$} sphere at step ${\bm{W}}_{i}$. Under weight constraint, projecting $\bm{G}$} onto the tangent space (Spectral Sphere) yields the largest update angle.
4 Algorithm Details
4.1 Spectral Radius Scale
Given the target spectral radius
$R=\Theta\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right)=c\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right)$
the constant $c$ serves as a scalar that sets the branch output magnitude relative to the residual stream. By tuning $c$, one can precisely control the signal-to-noise ratio along the deep residual path, balancing the contributions of the Attention/FFN blocks against the skip connection. Properly choosing $c$ is therefore essential for stabilizing depth-wise signal propagation in Transformers.
((a)) Radius scale search.
((b)) AbsMax of FFN activations.
((c)) RMS of FFN activations.
Figure 5: Ablation of radius scaling on optimization and activation. (a) Final loss vs. learning rate for varying radius scales $c$. A moderate scale (e.g. $c=2.0$) achieves the best performance. (b) AbsMax and (c) RMS of FFN activations during training. AbsMax monotonically follows the radius scale, whereas RMS follows a clear power-law scaling with $c$.
4.2 Learning Rate Scaler
We propose a unified view in which each learning-rate scaler enforces a consistent effective step size under a chosen norm metric and initialization scheme.
In the update rule ${\bm{W}}\leftarrow{\bm{W}}-\eta R{\bm{\Phi}}$, $R$ scales the update size. To avoid instability caused by heterogeneous layer shapes, we select $R$ to maintain a constant effective step size — defined as the ratio of update-to-weight magnitude under a norm metric $\|\cdot\|$ :
$\frac{\|\Delta{\bm{W}}\|}{\|{\bm{W}}\|}=\frac{\|\eta R{\bm{\Phi}}\|}{\|{\bm{W}}\|}\approx\eta.$ (15)
We evaluate three common learning rate scalers below:
$R=\begin{cases}\sqrt{d_{\mathrm{out}}/d_{\mathrm{in}}},&\text{({Spectral $\bm{\mu}$P Scaler)}}\\[6.0pt] \sqrt{\max(d_{\mathrm{out}},d_{\mathrm{in}})}\cdot 0.2,&\text{({Align-Adam-RMS Scaler})}\\[6.0pt] \sqrt{\max(1,d_{\mathrm{out}}/d_{\mathrm{in}})},&\text{({Spectral Kaiming Scaler})}\end{cases}$ (16)
Figure 6: Ablation of learning rate scalers. Each curve represents the optimal validation loss obtained via a learning rate grid search. Spectral $\bm{\mu}$P (yellow) outperforms Align-Adam-RMS (blue), validating the optimality of $\bm{\mu}$P-aligned scaling under the Spectral $\bm{\mu}$P condition.
  • Spectral $\bm{\mu}$P Scaler. Enforces the RMS-to-RMS operator norm invariance from Section˜2.1. It ensures that both $\bm{W}$} and $\Delta{\bm{W}}$ satisfy the $\bm{\mu}$P scaling $\|{\bm{W}}\|_{2}=\Theta(\sqrt{d_{\mathrm{out}}/d_{\mathrm{in}}})$, forming the geometric basis of our Spectral Sphere Optimizer.
  • Align-Adam-RMS Scaler. A heuristic for consistent relative learning rates in the RMS norm under fixed standard-deviation initialization. Empirically, it aligns per-layer update RMSnorm to AdamW, enabling direct transfer of AdamW-tuned hyperparameters (e.g. learning rate, weight decay) to the spectral method (liu2025muonscalablellmtraining).
  • Spectral Kaiming Scaler. Targets the spectral norm under Kaiming initialization (${\bm{W}}\sim\mathcal{N}(0,1/d_{\mathrm{in}})$(He2015Initialization). Random matrix theory establishes that such matrices satisfy $\|{\bm{W}}\|_{2}\approx 1+\sqrt{d_{\mathrm{out}}/d_{\mathrm{in}}}$. This scaling prevents vanishing pre-activations in bottleneck layers where $d_{\mathrm{out}}\ll d_{\mathrm{in}}$ (pethick2025trainingdeeplearningmodels).
Experimental results (Figure˜6) favor the Spectral $\bm{\mu}$P scaler, supporting the theoretical scaling condition in Section˜2.1. This demonstrates that optimizing on a spectral manifold requires a scaler explicitly calibrated to the spectral norm.
4.3 Module Granularity
Figure 7: Ablation of module granularity for initialization and optimization. Splitting QKV per head alone yields the most significant performance gain. Although splitting the FFN gate/up weights produces nearly the same loss curve as no-split, we maintain this split to respect their distinct functional roles; the overlapping curve is omitted from the figure.
To maximize computational efficiency, modern Transformer implementations such as Megatron-LM (megatron-lm) fuse several matrices into a single physical tensor (e.g. the QKV projections or the SwiGLU gate/up matrices). However, since these modules have distinct functional roles, imposing a unified constraint on the fused tensor is suboptimal as shown in Figure˜7. Instead, we decompose fused tensors and treat each submatrix as an independent module. We apply per module spectral initialization and optimization at this finer granularity by default (e.g. splitting attention QKV per head and separating FFN gate/up). Note that granularity is tunable depending on infra speed.
Under our spectral $\bm{\mu}$P initialization scheme (also discussed in Section˜4.2), the weight matrix is constructed by first sampling ${\bm{W}}_{k}\sim\mathcal{N}(0,\sigma^{2})$ and subsequently projecting it onto the spectral sphere:
${\bm{W}}=c\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{in}}}}\cdot\frac{{\bm{W}}_{k}}{\|{\bm{W}}_{k}\|_{2}},$
5 Infrastructure Design
5.1 Bottleneck Analysis
The main challenge in SSO implementation comes from the bracket-and-bisect root solver that runs at every update. To satisfy the tangent-space constraint, we must find a Lagrange multiplier $\lambda$ such that $h(\lambda)=0$. We first expand the search interval by bracketing, and then we run bisection inside the bracket to obtain a $\lambda^{\star}$ that meets a target tolerance. This solver introduces non-trivial overhead:
  1. 1.
    Workload Imbalance: different bracketing range and tolerance settings can change the number of search and bisection steps, leading to unstable runtime and workload imbalance between devices.
  2. 2.
    Computational Cost: Each step in the search also evaluates $h(\lambda)$, which calls $\operatorname{msign}(\widehat{{\bm{M}}}+\lambda{\bm{\Theta}})$; these extra matrix computations accumulate and add noticeable cost to every optimizer step.
  3. 3.
    Synchronization Overhead: Iterative search creates frequent synchronization between the GPU and the CPU, because the algorithm must finish each evaluation and then check a scalar condition before choosing the next $\lambda$ and continuing.
5.2 Optimization Pipeline
We introduce a holistic optimization pipeline designed to mitigate these overheads while preserving numerical precision. All performance statistics are collected from Dense 1.7B model pretraining.
Atomic Module Sharding.
We employ a fine-grained, parameter-wise sharding strategy (emergeing_optimizer). As noted by (liu2025muonscalablellmtraining), while standard ZeRO-1 is efficient for element-wise optimizers (e.g. AdamW), its flat-buffer sharding approach is incompatible with spectral methods that require full gradient matrices to compute updates. To reconcile this, we shard parameters as atomic modules rather than flattened buffers. An atomic module is defined as the minimal independent weight matrix required to remain intact for spectral operations (see Section˜4.3).
Load Balancing Strategy.
To resolve the workload imbalance caused by variable solver depths, we employ a “ping-pong” load balancing strategy adapted from (emergeing_optimizer). Empirical results indicate this method outperforms both greedy size-descent sorting and default round-robin allocation. We sort atomic modules by size and assign them to DP ranks in an alternating zigzag pattern. This heuristic effectively balancing the solver workload without complex runtime scheduling. Finally, we synchronize the updated params using iterative All-Gather collective from emergeing_optimizer.
Adaptive Kernel Selection.
We observe that the optimal implementation for Matrix Sign computation is highly sensitive to matrix dimensions. As shown in Table˜1, applying specialized kernels indiscriminately can degrade performance. We therefore implement an Adaptive Dispatcher:
  • Small Matrices (<512<512): We use a JIT-compiled PyTorch implementation built on torch.addmm. This avoids the launch overhead of specialized kernels.
  • Large Matrices ($\geq 512$): We dispatch to a custom Triton kernel implementing the SYmmetric Rank-K (SYRK) emergeing_optimizer update, which exploits the symmetry of Newton–Schulz iterations to halve memory reads.
Multi-Stream Parallelism.
For layers composed of many small independent matrices (e.g. per head attention split), single-stream execution suffers from kernel launch latency bubbles. We exploit this independence by dispatching spectral updates across multiple CUDA streams.
Mixed-Precision.
The Power Iteration for spectral norm estimation is performed in BFloat16, while the sensitive msign remains in FP32 with 8 iterations.
Cache Top Singular Vectors.
The singular vectors of model weights evolve slowly during training. Leveraging this temporal locality, we initialize the current Power Iteration using the cached singular vectors $u$ and $v$ from the previous step. This reuse mechanism drastically accelerates convergence, requiring only a few iterations to maintain approximation accuracy.
Table 1: Optimization breakdown on end-to-end latency for 4M tokens/step on NVIDIA B200. $\downarrow$ denotes improvement, while $\uparrow$ denotes regression. Note there is still room for improvement.
Time (ms) $\Delta$ vs Baseline $\Delta$ vs Prev.
Naive Baseline (No opt.) 10928.5 - -
+ Load balance & All Gather 9365.5 -1563.0 (-14.3%) $\downarrow$ -1563.0 (-14.3%) $\downarrow$
+ Triton SYRK Kernel 10284.2 -644.3  (-5.9%) $\downarrow$ +918.7  (+9.8%) $\uparrow$
+ Adaptive & Multi-stream 9383.4 -1545.1 (-14.2%) $\downarrow$ -900.8  (-8.8%) $\downarrow$
+ BF16 & Torch.compile (Final) 7666.3 -3262.2 (-29.9%) $\downarrow$ -1717.1 (-18.3%) $\downarrow$
Table 2: End-to-end per-step latency for 4M tokens/step on NVIDIA B200. Muon as baseline.
AdamW Muon MuonSphere Spectral Sphere
Time (ms) 6734.15 (-2.10%) 6878.83 6949.85 (+1.03%) 7666.32 (+11.45%)
5.3 Future Improvements
  • GPU-Native Solver. Profiling indicates that the current CPU-based bisection solver introduces latency due to frequent device-host synchronization. Although the bracketing phase converges rapidly (<2<2 steps), the bisection phase averages 5–7 steps, creating synchronization bubbles. Future work will prioritize implementing a pure GPU-native solver to eliminate these overheads. Additionally, we plan to adopt higher convergence algorithms, such as Brent’s method or $n$-section search, and optimize initial bracketing intervals to minimize msign calls.
  • Kernel Optimization. The theoretical advantage of SSO relies on the accuracy of the tangent space projection; when errors are high, the update direction may degrade to the “worst-case” Muon update. Currently, we ensure precision via 8 iterations of msign in FP32. To follow standard Muon practices, we plan to use msign in BFloat16 within 5 steps. Furthermore, we intend to replace current JIT-compiled operations with custom, fully optimized kernels (e.g. for batched msign and Power Iterations) to better exploit the hardware features of next-generation GPUs.
  • Low-precision Training. While weight matrics are spectral constrained, we found the residual stream remains the primary source of outliers. We aim to explore “fully manifold constrained” architectures, i.e. using mHC (xie2025mhcmanifoldconstrainedhyperconnections). Additionally, given SSO’s demonstrated stability, we plan to explore low-precision training (e.g. FP8/NVFP4) to leverage the high throughput of low-bit arithmetic for better training efficiency.
6 Scaling Experiments
6.1 Experimental Setup
Hyperparameters.
Following the $\bm{\mu}$P protocol, we perform a learning rate sweep on the 1.7B model with AdamW between $10^{-3}$, $10^{-2}$] and found $5\times 10^{-3}$ to be the optimal value. Training uses 500 warmup steps, a global batch size of $\sim$4M tokens (1024 sequences $\times$ 4096 tokens each), and a cosine learning rate decay reduced to $10\%$ peak. We use BF16 mixed-precision training.
Following Kimi Moonlight (liu2025muonscalablellmtraining), all optimizers use a weight decay of 0.1. However, distinct from their approach of aligning updates to AdamW RMS (intended to reuse current scaling laws), we find that spectral $\bm{\mu}$P LR scaler outperforms the uniform 0.2 update-rms alignment (Section˜4.2). For msign coefficients, we use the Polar Express method (amsel2025polarexpressoptimalmatrix) with 8 Newton–Schulz iterations [^6] We tested 5 and 8 iterations as well as different msign coefficients and observed negligible differences in training loss (¡ 1e-3). We retain 8 iterations for improved numerical accuracy.. We employ Nesterov momentum and, by default, split attention heads and FFN gate/up projections for separate initialization and optimization. For Spectral Sphere, we set the $\lambda$-solver’s maximum iterations to 20, Lagrange solver precision tolerance to 2e-4, and remove weight decay for all hidden 2D weights, as retraction maintains the weight constraint (Section˜3.3).
Training Data
We use the OLMo-Mix-1124 dataset (olmo20252olmo2furious), tokenized with the OLMo-2 tokenizer. We randomly sample 100 billion tokens for training and reserve 1 billion tokens for validation. Data indices are built offline to guarantee a deterministic training order.
Benchmarks.
We primarily evaluate on the following downstream tasks: ARC (clark2018arc), BoolQ (clark2019boolq), CSQA (talmor2019csqa), HellaSwag (zellers2019hellaswag), PIQA (bisk2020piqa), WinoGrande (sakaguchi2021winogrande) and LAMBADA (paperno2016lambada).
6.2 Dense 1.7B
We adopt the architecture configuration of Qwen3-1.7B (yang2025qwen3technicalreport), replacing its original tokenizer with that of OLMo-2 (olmo20252olmo2furious). The core architecture utilizes Grouped Query Attention (GQA), QK-Norm, SwiGLU activations, Rotary Positional Embeddings (RoPE), and pre-normalization RMSNorm. We do not tie the embedding and the head.
Figure 8: Validation loss of training dense 1.7B model on 100B tokens. As a reference point, AdamW attains a final validation loss of 2.588 at 23k steps. The overall setup favors AdamW, since the learning rate is set for it (5e-3), rather than the higher optimal rate (1e-2) for Muon and Spectral Sphere. Even under this setting, spectral-based optimizers exhibit higher efficiency: Muon reaches the same validation loss level in 12% fewer steps, while Spectral Sphere does so in 19% fewer steps.
Table 3: Performance comparison of different optimizers on Dense 1.7B models.
Optimizer LMB. LMB. CSQA PIQA Hella. Wino. ARC-e ARC-c BoolQ Avg.
(PPL) $\downarrow$ (Acc) $\uparrow$ (Acc) $\uparrow$ (Acc) $\uparrow$ (Acc) $\uparrow$ (Acc) $\uparrow$ (Acc) $\uparrow$ (Acc) $\uparrow$ (Acc) $\uparrow$ (Acc) $\uparrow$
AdamW 5.40 63.71 19.66 74.70 47.90 62.59 68.81 37.37 63.24 54.75
Muon 5.05 65.19 19.00 75.35 48.91 61.72 70.24 37.46 64.22 55.26
MuonSphere 4.87 65.55 20.07 74.97 49.20 62.83 71.51 38.40 66.97 56.19
Spectral Sphere 5.00 65.07 21.05 75.95 49.25 63.77 71.80 38.31 65.57 56.35
6.3 MoE 8B-A1B
The configuration largely follows DeepSeek-V3 (deepseekai2025deepseekv3technicalreport). The model has 27 layers: the first is a standard dense FFN, followed by 26 MoE layers. We utilize 64 experts in total, with a top-4 routing expert plus 1 shared expert.
Router and Load Balancing. The router is implemented in FP32 precision. We adopt the auxiliary-loss-free strategy (wang2024auxiliarylossfreeloadbalancingstrategy), which demonstrated superior expert loading balance compared to global auxiliary loss (qiu2025demonsdetailimplementingload) in our ablations. We enable expert bias with an update rate of 0.001. The sequence-level auxiliary loss is removed, as we found it redundant when expert bias is enabled. we use a sigmoid gate with top-$k$ scores renormalized and scaled by 2. (see Section˜A.4). To evaluate router load balancing, we employ Maximal Violation (MaxVio) (wang2024auxiliarylossfreeloadbalancingstrategy), where 0 indicates perfect balance. As shown in Figure˜9, weight spectral normalization effectively promotes balanced routing, leading to superior validation loss.
((a)) Validation Loss across four optimizers.
((b)) Max Vio as a MoE load-balance metric.
Figure 9: MoE 8B-A1B training. Spectral Sphere achieves the lowest validation loss while maintaining the best expert load balance. In contrast, AdamW exhibits substantially larger MaxVio with frequent spikes, indicating unstable routing and poorer utilization of effective model capacity. Compared to Muon, constraining each expert on the spectral sphere further improves load balance.
6.4 DeepNet 200-Layer
To evaluate the stability of different optimizers under extreme depth, we extended the baseline’s 28 layers to 200 layers. This deep and narrow serves as a stress test for stability. The training loss in Figure˜10 shows that Spectral Sphere outperforms baselines with lower loss and higher stability.
Figure 10: Deepnet 200 layers training loss. AdamW shows pronounced instability, characterized by frequent loss spikes and a growing gap in performance relative to spectral-based optimizers. Spectral Sphere attains both the lowest loss and the highest stability.
7 Discussion
In this work, we propose a novel perspective on optimizer design grounded in spectral $\bm{\mu}$P principles. As detailed in Section˜2.1, $\bm{\mu}$P provides a mathematical safeguard that strictly controls hidden layer activations at the desired scale. By identifying the spectral sphere as the natural geometry for stable feature learning, we derive the Spectral Sphere Optimizer (SSO) — the unique solution for steepest descent constrained within both weight and update manifolds (Section˜3). This formulation effectively achieves rapid convergence grounded in fundamental training stability. Empirically, SSO consistently outperforms AdamW and Muon, while uniquely preserving stable $\bm{\mu}$P learning rate transfer. Notably, it yields superior training dynamics: it significantly improves MoE router load balancing, suppresses outliers in deep networks, and strictly bounds activations within a tunable scale.
A critical distinction exists between our approach and emerging works on Stiefel manifold optimization (bernstein2025manifolds): while Stiefel manifold requires all singular values to be exactly 1, SSO constrains only the maximal singular value. This relaxation allows the internal spectrum to evolve freely below the bound, avoiding the overly rigid isotropy of the Stiefel manifold.
Beyond the theoretical contribution, we provide a complete and systematic recipe (Section˜4) implemented in Megatron-LM. Our guidelines on atomic granularity, “ping-pong” load balancing strategy, learning rate scaler, and spectral radius scale serve as a robust empirical practice for the broader class of spectral optimizers. Finally, while we acknowledge that the current root solver introduces non-trivial latency, we have outlined concrete pathways to mitigate this overhead in Section˜5.3. For scenarios prioritizing infra cost, we recommend MuonSphere — a variant that retains equivalent activation control with minimal overhead.
Acknowledgements
We sincerely thank Jianlin Su for his excellent blog, Scientific Spaces.
Appendix A Appendix
A.1 Duality with Spectral Norm
We may first prove Theorem˜A.1, and we can then directly derive Equation˜11.
Theorem A.1.
Nuclear norm $\|\!\cdot\!\|_{*}$ is the dual norm of spectral norm $\|\!\cdot\!\|_{2}$, which means
$\|{\bm{G}}\|_{*}=\max_{\|T\|_{2}=1}\langle{\bm{G}},{\bm{T}}\rangle.$
$\operatorname{msign}(\cdot)$ is the dual map based on $\|\!\cdot\!\|_{2}$, which means
$\operatorname{msign}({\bm{G}})=\underset{\|{\bm{T}}\|_{2}=1}{\operatorname*{argmax}}\langle{\bm{G}},{\bm{T}}\rangle.$
Proof.
Since ${\bm{G}}\in{\mathbb{R}}^{n\times m}$ has Singular Vector Decomposition ${\bm{G}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{\top}=\sum_{i=1}^{r}{\sigma}_{i}{\bm{u}}_{i}{\bm{v}}_{i}^{\top}$ where ${\bm{u}}_{i}\in{\mathbb{R}}^{n}$ and ${\bm{v}}_{i}\in{\mathbb{R}}^{m}$ are left and right singular vectors, we have
$\langle{\bm{G}},{\bm{T}}\rangle=\operatorname{tr}({\bm{G}}^{\top}{\bm{T}})=\operatorname{tr}(\sum_{i=1}^{r}{\sigma}_{i}{\bm{v}}_{i}{\bm{u}}_{i}^{\top}{\bm{T}})=\sum_{i=1}^{r}{\sigma}_{i}{\bm{u}}_{i}^{\top}{\bm{T}}{\bm{v}}_{i}.$ (17)
With $\|{\bm{T}}\|_{2}=1$ and $\|{\bm{u}}_{i}\|_{2}=\|{\bm{v}}_{i}\|_{2}=1$, we have ${\bm{u}}_{i}^{\top}{\bm{T}}{\bm{v}}_{i}\leq\|{\bm{T}}\|_{2}\|{\bm{u}}_{i}\|_{2}\|{\bm{v}}_{i}\|_{2}=1$, hence
$\langle{\bm{G}},{\bm{T}}\rangle=\sum_{i=1}^{r}{\sigma}_{i}{\bm{u}}_{i}^{\top}{\bm{T}}{\bm{v}}_{i}\leq\sum_{i=1}^{r}{\sigma}_{i}=\|{\bm{G}}\|_{*}.$ (18)
Equality is attained when ${\bm{u}}_{i}^{\top}{\bm{T}}{\bm{v}}_{i}=1$ for all $i=1,\dots,r$, that is when
${\bm{T}}=\sum_{i=1}^{r}{\bm{u}}_{i}{\bm{v}}_{i}^{\top}={\bm{U}}_{[:,:r]}{\bm{V}}_{[:,:r]}^{\top}=\operatorname{msign}({\bm{G}}).$ (19)
Note that for this $\bm{T}$} we indeed have $\|{\bm{T}}\|_{2}=1$ (its nonzero singular values are all equal to $1$), so $\bm{T}$} is feasible and achieves the upper bound. Therefore, we have
$\underset{\|{\bm{T}}\|_{2}=1}{\operatorname*{argmax}}\langle{\bm{G}},{\bm{T}}\rangle=\operatorname{msign}({\bm{G}}),$ (20)
and according to Equation˜18, the maximum value equals $\|{\bm{G}}\|_{*}$. ∎
A.2 Proofs: Localization of the Root of $\bm{h(\lambda)}$
In this section, we first prove the localization of the root $\lambda^{\star}$ to $h(\lambda)=0$, which is required in Section˜3.2. We then present experiments showing that the computational overhead of the proposed $\lambda$-solver is negligible compared to the whole training process.
We first prove Theorem˜A.2, from which Theorem˜A.3 follows.
Theorem A.2.
The function
$h(\lambda)=\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda)\rangle=\langle{\bm{\Theta}},\operatorname{msign}({\bm{G}}+\lambda{\bm{\Theta}})\rangle$
is monotonic non-decreasing in $\lambda$.
Proof.
Recall that the matrix sign function can be equivalently defined as the solution of the spectral-norm constrained maximization as in Theorem˜A.1:
$\operatorname{msign}({\bm{G}})=\underset{\|T\|_{2}=1}{\operatorname*{argmax}}\langle{\bm{G}},{\bm{T}}\rangle.$ (21)
Consider two values $\lambda_{2}>\lambda_{1}$. By the definition of ${\bm{\Phi}}^{\star}(\lambda)$ as the maximizer of $\langle{\bm{G}}+\lambda{\bm{\Theta}},\cdot\rangle$, we obtain
$\langle{\bm{G}}+\lambda_{1}{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle\geq$ $\langle{\bm{G}}+\lambda_{1}{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle$
$\Rightarrow\quad$ $\langle{\bm{G}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle+\lambda_{1}\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle\geq\langle{\bm{G}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle+\lambda_{1}\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle,$ (22)
$\langle{\bm{G}}+\lambda_{2}{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle\geq$ $\langle{\bm{G}}+\lambda_{2}{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle$
$\Rightarrow\quad$ $\langle{\bm{G}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle+\lambda_{2}\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle\geq\langle{\bm{G}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle+\lambda_{2}\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle.$ (23)
Summing Equation˜22 and Equation˜23, we can derive
$\lambda_{1}\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle+\lambda_{2}\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle\geq\lambda_{1}\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle$ $+\lambda_{2}\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle$ (24)
$\Rightarrow\quad(\lambda_{2}-\lambda_{1})\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle\geq(\lambda_{2}-\lambda_{1})\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle\quad$ $\Rightarrow\quad\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{2})\rangle\geq\langle{\bm{\Theta}},{\bm{\Phi}}^{\star}(\lambda_{1})\rangle,$ (25)
with $\lambda_{2}-\lambda_{1}>0$. Hence, $h(\lambda)=\langle{\bm{\Theta}},\operatorname{msign}({\bm{G}}+\lambda{\bm{\Theta}})\rangle$ is monotonic non-decreasing in $\lambda$. ∎
Theorem A.3.
There exists some $\lambda^{\star}\in{\mathbb{R}}$ such that $h(\lambda^{\star})=0$, and any such root satisfies
$|\lambda^{\star}|\leq 2\|{\bm{G}}\|_{*}.$
Proof.
We first prove the existence of a root $\lambda^{\star}$. Since ${\bm{\Theta}}={\bm{u}}_{1}{\bm{v}}_{1}^{\top}$ is a rank-one matrix with unit-norm singular vectors, its nuclear norm satisfies $\|{\bm{\Theta}}\|_{*}=1$. Moreover, by construction, we have
$\|{\bm{\Phi}}^{\star}(\lambda)\|_{2}=\|\operatorname{msign}({\bm{G}}+\lambda{\bm{\Theta}})\|_{2}=1.$ (26)
By Theorem˜A.1, letting ${\bm{T}}_{1}=\frac{{\bm{\Phi}}^{\star}(\lambda)}{\|{\bm{\Phi}}^{\star}(\lambda)\|_{2}},{\bm{T}}_{2}=-\frac{{\bm{\Phi}}^{\star}(\lambda)}{\|{\bm{\Phi}}^{\star}(\lambda)\|_{2}}$, we have
$\|{\bm{T}}_{1}\|_{2}=\|{\bm{T}}_{2}\|_{2}=1\quad\text{and thus}\quad\langle{\bm{\Theta}},{\bm{T}}_{1}\rangle\leq\|{\bm{\Theta}}\|_{*}\ \text{and}\ \langle{\bm{\Theta}},{\bm{T}}_{2}\rangle\leq\|{\bm{\Theta}}\|_{*}.$ (27)
Therefore, we have
$\left|\langle{\bm{\Theta}},\operatorname{msign}({\bm{G}}+\lambda{\bm{\Theta}})\rangle\right|\leq\|{\bm{\Theta}}\|_{*}\,\|\operatorname{msign}({\bm{G}}+\lambda{\bm{\Theta}})\|_{2}=1,$ (28)
which implies $|h(\lambda)|\leq 1$ for all $\lambda\in{\mathbb{R}}$. Moreover, since $h(\lambda)$ is monotonic non-decreasing by Theorem˜A.2, the limits $\lim_{\lambda\to-\infty}h(\lambda)$ and $\lim_{\lambda\to+\infty}h(\lambda)$ exist. Using the property $\operatorname{msign}(\lambda{\bm{T}})=\operatorname{msign}({\bm{T}})$, we have
$\lim_{\lambda\to+\infty}\operatorname{msign}({\bm{G}}+\lambda{\bm{\Theta}})=\lim_{\lambda\to+\infty}\operatorname{msign}\left[\lambda\left({\bm{\Theta}}+\tfrac{1}{\lambda}{\bm{G}}\right)\right]=\operatorname{msign}({\bm{\Theta}})={\bm{\Theta}},$ (29)
where the last equality is satisfied by definition of $\operatorname{msign}(\cdot)$ and $\bm{\Theta}$}. Consequently,
$\lim_{\lambda\to+\infty}h(\lambda)=\langle{\bm{\Theta}},{\bm{\Theta}}\rangle=\operatorname{tr}({\bm{v}}_{1}{\bm{u}}_{1}^{\top}{\bm{u}}_{1}{\bm{v}}_{1}^{\top})=1.$ (30)
Similarly, we can also obtain $\lim_{\lambda\to-\infty}h(\lambda)=-1$. Since $\operatorname{msign}({\bm{G}}+\lambda{\bm{\Theta}})$ is a subgradient of a convex function $\|{\bm{G}}+\lambda{\bm{\Theta}}\|_{*}$ with respect to $\lambda$ according to watson1992characterization, monotonic non-decreasing $h(\lambda)$ satisfies the intermediate value property. Consequently, there exists at least one $\lambda^{\star}\in{\mathbb{R}}$ such that $h(\lambda^{\star})=0$.
We next localize the root. For any $\lambda>2\|{\bm{G}}\|_{*}>0$ and any matrix $\bm{T}$} with $\|{\bm{T}}\|_{2}=1$, we have
$\lambda\langle{\bm{\Theta}},{\bm{T}}\rangle=\langle{\bm{G}}+\lambda{\bm{\Theta}},{\bm{T}}\rangle-\langle{\bm{G}},{\bm{T}}\rangle.$ (31)
By Equation˜27, $\langle{\bm{G}},{\bm{T}}\rangle\leq\|{\bm{G}}\|_{*}$. Let ${\bm{T}}=\operatorname{msign}({\bm{G}}+\lambda{\bm{\Theta}})$, using Theorem˜A.1, we can derive
$\lambda h(\lambda)=\|{\bm{G}}+\lambda{\bm{\Theta}}\|_{*}-\langle{\bm{G}},{\bm{T}}\rangle\geq\big|\|{\bm{G}}\|_{*}-\lambda\|{\bm{\Theta}}\|_{*}\big|-\|{\bm{G}}\|_{*}\geq\lambda-2\|{\bm{G}}\|_{*}>0,$ (32)
[…]
Theorem˜A.3 actually provides a theoretical guarantee that the correctness of using the bisection algorithm to solve the root $\lambda^{\star}$. To further validate this theoretical result, we randomly generate several zero-centered matrices with varying dimensions, construct $\bm{G}$} and $\bm{\Theta}$} as described in this section, and plot the curve of $h(\lambda)$ together with its root. As shown in Figure˜3, $f(\lambda)$ is indeed monotonic in $\lambda$, and its root lies close to $\lambda=0$, which is consistent with our theoretical analysis.
A.3 Dynamic Spectral Weight Decay
Figure 11: Abaltion of weight decay in Muon. In contrast with Spectral Sphere (Figure˜1), without weight constraint, Muon training dynamics become instable, which in-turn would hurt performance.
In this section, we introduce spectral retraction mechanisms to counteract the accumulation of higher-order errors that may gradually drift the weights off the spectral sphere. As noted in Section˜3.3, the remainder term $\mathcal{O}(\eta^{2}R^{2}\|{\bm{\Phi}}\|_{2}^{2})$ in the Taylor expansion can accumulate over iterations. To enforce the exact constraint $\|{\bm{W}}\|_{2}=R$ throughout training, we apply Equation˜13 that projects the weights back onto the spectral manifold.
In practice, we consider two retraction variants that differ in how strictly the spectral constraint is enforced. The hard variant applies an explicit projection to maintain $\|{\bm{W}}\|_{2}=R$ at every step, while the dynamic variant replaces the exact projection with a soft, learning-rate-scaled radial correction that steers $\|{\bm{W}}\|_{2}$ toward the target radius. These two variants correspond to enforcing the spectral constraint exactly or approximately, and lead to different interactions with conventional weight decay.
Hard retraction (exact projection).
The hard variant applies the retraction map explicitly using an estimated top singular value $\sigma\approx\|{\bm{W}}\|_{2}$ which is computed via Power Iteration.
${\bm{W}}\leftarrow\frac{R}{\sigma}\,{\bm{W}}.$ (33)
This enforces the exact constraint $\|{\bm{W}}\|_{2}=R$ at every step, with deviations arising solely from the approximation error in $\sigma$. Under this setting, Spectral Sphere is typically used together with standard decoupled weight decay as in AdamW: weight decay shrinks the full parameter vector, while the spectral retraction constrains the dominant singular value.
Dynamic retraction (soft spectral decay).
The dynamic variant replaces the exact projection with a small, sign-controlled radial adjustment with hyperparameter $\lambda$.
${\bm{W}}\leftarrow\bigl(1+\lambda\,\eta\operatorname{sign}(R-\sigma)\bigr){\bm{W}}.$ (34)
Instead of strictly enforcing the constraint $\|{\bm{W}}\|_{2}=R$, this update gently adjusts the spectral norm toward the target radius $R$ based on $\sigma$. Since the correction is proportional to $\eta$, the adjustment strength diminishes automatically as learning rate schedules. Under this formulation, dynamic retraction functions as a spectrally aligned, layer-wise analogue to decoupled weight decay. Consequently, we tie $\lambda$ to the AdamW weight decay coefficient and employ the dynamic variant as a drop-in replacement for conventional weight decay.
From a geometric perspective, the two variants above can both be viewed as approximate projections onto the spectral sphere. As discussed in Section˜3.3, applying retraction at the beginning of iteration $t+1$ instead of after the update at iteration $t$ only introduces $\mathcal{O}(\eta^{2})$ discrepancies. Consequently, the update remains first-order equivalent to steepest descent on the spectral manifold.
A.4 MoE Routing Scaling Factor
Following  [kexuefm-10945], in our MoE architecture, the output $\bm{y}$} is the sum of the shared experts and the routed experts. A critical issue arises from the magnitude mismatch between the two parts:
${\bm{y}}=\underbrace{\sum_{i=1}^{N_{\text{shared}}}{\bm{e}}_{\text{shared},i}}_{\text{Weight }\approx 1}+\underbrace{\sum_{j\in\text{TopK}}g_{j}{\bm{e}}_{\text{routed},j}}_{\text{Weight }g_{j}\ll 1}$ (35)
The shared experts are directly added to the residual stream, effectively having a coefficient of 1. In contrast, the routed experts are multiplied by sigmoid probabilities $g_{j}$, which are typically small. Consequently, the variance (signal energy) of the routed experts is significantly lower than that of the shared experts, causing the optimizer to neglect the routed part.
To balance the contributions, we compute a scaling factor $M$ to the routed experts:
${\bm{y}}=\sum_{i=1}^{N_{\text{shared}}}{\bm{e}}_{\text{shared},i}+M\sum_{j\in\text{TopK}}g_{j}{\bm{e}}_{\text{routed},j}$ (36)
We aim to choose $M$ such that the expected variance of the routed term matches that of the shared term, under assumption that experts have unit variance:
$M\approx\sqrt{\frac{N_{\text{shared}}}{\mathbb{E}[\sum g_{j}^{2}]}}$ (37)
Using numerical simulation (LABEL:lst:moe_scaling) with $N_{\text{shared}}=1$ and Top-4 sigmoid routing, we find $M\approx 2.0$. We find this scaling factor is crucial for MoE training stability.
Listing 1: Python implementation for estimating the MoE scaling factor $M$.
1import numpy as np
2
3def estimate_scaling_factor(n_total=64, k_routed=4, n_shared=1):
4 factors = []
5 for _ in range(10000):
6 # 1. Simulate logits
7 logits = np.random.randn(n_total - n_shared)
8
9 # 2. Get Sigmoid Scores
10 scores = 1 / (1 + np.exp(-logits))
11
12 # 3. Select Top-K routed scores
13 scores = np.sort(scores)[::-1][:k_routed]
14
15 # 3. Normalize scores
16 scores /= scores.sum()
17
18 # 4. Calculate Magnitude
19 magnitude = np.sum(scores**2)**0.5
20
21 factors.append(n_shared**0.5 / magnitude)
22
23 return np.mean(factors)
A.5 $\bm{\mu}$P Width Scaling
In Figure˜2, we conduct experiment to validate $\bm{\mu}$P width scaling across different model sizes from 70M to 1.8B. Additional AdamW results are included in Figure˜12. We scale the hidden size, intermediate size, and number of attention heads (while fixing head dimensions). The models are trained on 30B tokens sampled from the Olmo2-mix124 dataset [olmo20252olmo2furious]. All optimizers use the Spectral $\bm{\mu}$P LR Scaler.
  • The LR is swept across {1e-3, 3e-3, 5e-3, 7e-3, 9e-3, 1e-2, 1.5e-2, 2e-2, 3e-2}.
  • Hidden size is swept across {256, 512, 1024, 2048}.
Figure 12: $\bm{\mu}$P LR grid search with AdamW, Muon and Spectral Sphere. AdamW shows even worse validation loss, with drifting optimal LR.

### v2

@@
 $\bm{\mu}$P scale invariance requires maintaining $\|{\bm{y}}\|_{\mathrm{rms}}=\|{\bm{x}}\|_{\mathrm{rms}}=\Theta(1),$ which is equivalent to enforcing the RMS-to-RMS condition $\|{\bm{W}}\|_{\mathrm{rms}\to\mathrm{rms}}=\|{\bm{W}}\|_{2}\sqrt{{d_{\mathrm{in}}}/{d_{\mathrm{out}}}}=\Theta(1).$ This induces the following spectral norm constraint on the weight matrix $\|{\bm{W}}\|_{2}=\Theta(\sqrt{{d_{\mathrm{out}}}/{d_{\mathrm{in}}}}).$ A similar requirement applies to parameter updates; together, we refer to these as the spectral $\bm{\mu}$P condition below.
-yang2023spectral shows that preserving scale-invariant activations for feature learning requires the same law to hold for both weights and their updates: $\|{\bm{W}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{int}}}}\right),\qquad\|{\bm{\Phi}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{int}}}}\right).$
+yang2023spectral shows that preserving scale-invariant activations for feature learning requires the same law to hold for both weights and their updates: $\|{\bm{W}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{in}}}}\right),\qquad\|{\bm{\Phi}}\|_{2}=\Theta\left(\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{in}}}}\right).$
 2.2 Steepest Descent under Different Norms
-Following metrized deep learning (bernstein2024oldoptimizernewnorm), many successful optimizers (e.g.
-AdamW (loshchilov2019decoupled), Shampoo (gupta2018shampoo), Prodigy (mishchenko2024prodigy)) can be interpreted as first-order methods without convexity assumptions.
-Specifically, after switching off exponential moving averages, these algorithms reduce to instances of steepest descent governed by distinct norms.
-Under this framework, an optimizer is fundamentally defined by a weight update $\bm{\Phi}$} that minimizes a quadratic model of the loss:
-${\bm{\Phi}}:=\underset{{\bm{\Phi}}}{\operatorname*{argmin}}\left\{\mathcal{L}({\bm{W}})+\langle{\bm{G}},{\bm{\Phi}}\rangle+\frac{\lambda}{2}\|{\bm{\Phi}}\|^{2}\right\}.$ (3)
-The update is thus determined by two priors: a norm $\|\!\cdot\!\|$ assigned according to the specific functional role of the module, and a sharpness $\lambda$ that governs the update scale.
+Following bernstein2024oldoptimizernewnorm; cesista2025sdcbs, many successful optimizers can be interpreted as first-order methods without convexity assumptions.
+Specifically, after switching off exponential moving averages, these algorithms reduce to instances of steepest descent governed by distinct norms: SGD corresponds to the Frobenius norm $\|\!\cdot\!\|_{F}$, AdamW (loshchilov2019decoupled) to the $\ell_{\infty}$ norm, and Shampoo (gupta2018shampoo) to the Spectral norm $\|\!\cdot\!\|_{2}$.
+In this view, we determine the update $\bm{\Phi}$} by iteratively minimizing a first-order Taylor approximation of the loss around the current weights $\bm{W}$}, subject to a hard constraint on the step size:
+${\bm{\Phi}}:=\underset{\|{\bm{\Phi}}\|\leq\eta}{\operatorname*{argmin}}\left\{\mathcal{L}({\bm{W}})+\left\langle\nabla_{\bm{W}}\mathcal{L}({\bm{W}}),{\bm{\Phi}}\right\rangle\right\}=\underset{\|{\bm{\Phi}}\|\leq\eta}{\operatorname*{argmin}}\langle{\bm{G}},{\bm{\Phi}}\rangle,$ (3)
+where ${\bm{G}}:=\nabla_{\bm{W}}\mathcal{L}({\bm{W}})$ is the gradient of the loss function at $\bm{W}$}.
+The update is thus determined by two priors: a norm $\|\!\cdot\!\|$ assigned according to the specific functional role of the module, and a step size $\eta$ that governs the update scale.
 The solution is as follows:
-Given a norm $\|\!\cdot\!\|$ that endows the parameter space with a geometry, the steepest descent update is ${\bm{\Phi}}=-\eta\cdot{\bm{\Phi}}_{\text{unit}},\quad\text{where}\quad{\bm{\Phi}}_{\text{unit}}:=\underset{\|{\bm{T}}\|\leq 1}{\arg\max}\langle{\bm{G}},{\bm{T}}\rangle,\quad\text{and}\quad\eta:=\frac{\|{\bm{G}}\|_{\dagger}}{\lambda}.$ (4) Here, $\|{\bm{G}}\|_{\dagger}:=\max_{\|{\bm{T}}\|\leq 1}\langle{\bm{G}},{\bm{T}}\rangle$ is the dual norm of the gradient $\bm{G}$} induced by $\|\!\cdot\!\|$.
-This perspective unifies diverse algorithms by mapping them to their underlying geometries: SGD corresponds to Frobenius norm $\|\!\cdot\!\|_{F}$, Adam to $l_{\infty}$ norm, and Shampoo to spectral norm $\|\!\cdot\!\|_{2}$.
+Given a norm $\|\!\cdot\!\|$ defining the geometry and a step size $\eta$ governing the scale, the steepest descent update solution to Equation˜3 is given by ${\bm{\Phi}}=-\eta\cdot{\bm{\Phi}}_{\text{unit}},\quad\text{where}\quad{\bm{\Phi}}_{\text{unit}}:=\underset{\|{\bm{T}}\|\leq 1}{\arg\max}\langle{\bm{G}},{\bm{T}}\rangle,$ (4)
 2.3 Muon Optimizer
@@
 […]
-$\max_{{\bm{\Phi}}}\ \langle{\bm{G}},{\bm{\Phi}}\rangle\quad\text{s.t.}\quad\|{\bm{\Phi}}\|_{2}\ =1,\ \ \langle{\bm{\Theta}},{\bm{\Phi}}\rangle=0.$ (10)
-We then introduce a Lagrange multiplier $\lambda$ and maximize Lagrangian $\mathcal{L}({\bm{\Phi}},\lambda)=\langle{\bm{G}}+\lambda{\bm{\Theta}},{\bm{\Phi}}\rangle$ under constraint $\|{\bm{\Phi}}\|_{2}=1$.
+$\max_{{\bm{\Phi}}}\ \langle{\bm{G}},{\bm{\Phi}}\rangle\quad\text{s.t.}\quad\|{\bm{\Phi}}\|_{2}\ =1,\ \ \langle{\bm{\Theta}},{\bm{\Phi}}\rangle=0.eq$ (10)
+We then introduce a Lagrange multiplier $\lambda$ and maximize Lagrangian $\mathcal{L}({\bm{\Phi}};\lambda)=\langle{\bm{G}}+\lambda{\bm{\Theta}},{\bm{\Phi}}\rangle$ under constraint $\|{\bm{\Phi}}\|_{2}=1$.
 The analytical solution and numerical method are summarized below.
@@
 As a result, weight decay, which is primarily introduced to limit the weight scale, becomes redundant.
-We therefore eliminate weight decay in hidden 2D weights [^4] We still apply weight decay to 1D params (e.g.
-Embedding, RMSNorm) for possible scaling stability, although our ablations on 1.7B model suggest that disabling 1D params may actually be slightly better., removing a sensitive hyperparameter from training.
+We therefore eliminate weight decay in hidden 2D weights [^4] We still apply weight decay to params like Embedding for possible scaling stability, although our ablations on 1.7B model suggest that disabling weight decay may actually be slightly better., removing a sensitive hyperparameter from training.
 More details can be found in Section˜A.3.
@@
 […]
-[…]
 Figure 4: Geometry of Steepest Descent Update Directions.
@@
 Given the target spectral radius
-$R=\Theta\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right)=c\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right)$
+$R=\Theta\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right)=c\left(\sqrt{d_{\text{out}}/d_{\text{in}}}\right),$
 the constant $c$ serves as a scalar that sets the branch output magnitude relative to the residual stream.
@@
 4.2 Learning Rate Scaler
-We propose a unified view in which each learning-rate scaler enforces a consistent effective step size under a chosen norm metric and initialization scheme.
+We propose a unified view in which each learning rate scaler enforces a consistent effective step size under a chosen norm metric and initialization scheme.
 In the update rule ${\bm{W}}\leftarrow{\bm{W}}-\eta R{\bm{\Phi}}$, $R$ scales the update size.
@@
 Under our spectral $\bm{\mu}$P initialization scheme (also discussed in Section˜4.2), the weight matrix is constructed by first sampling ${\bm{W}}_{k}\sim\mathcal{N}(0,\sigma^{2})$ and subsequently projecting it onto the spectral sphere:
-${\bm{W}}=c\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{in}}}}\cdot\frac{{\bm{W}}_{k}}{\|{\bm{W}}_{k}\|_{2}},$
+${\bm{W}}=c\sqrt{\frac{d_{\mathrm{out}}}{d_{\mathrm{in}}}}\cdot\frac{{\bm{W}}_{k}}{\|{\bm{W}}_{k}\|_{2}}.$
 5 Infrastructure Design
@@
   • Low-precision Training.
-While weight matrics are spectral constrained, we found the residual stream remains the primary source of outliers.
+While weight matrics are spectral constrained, we find the residual stream remains the primary source of outliers.
 We aim to explore “fully manifold constrained” architectures, i.e.
@@
 Hyperparameters.
-Following the $\bm{\mu}$P protocol, we perform a learning rate sweep on the 1.7B model with AdamW between $10^{-3}$, $10^{-2}$] and found $5\times 10^{-3}$ to be the optimal value.
+Following the $\bm{\mu}$P protocol, we perform a learning rate sweep on the 1.7B model with AdamW between $10^{-3}$, $10^{-2}$] and find $5\times 10^{-3}$ to be the optimal value.
 Training uses 500 warmup steps, a global batch size of $\sim$4M tokens (1024 sequences $\times$ 4096 tokens each), and a cosine learning rate decay reduced to $10\%$ peak.
@@
 For Spectral Sphere, we set the $\lambda$-solver’s maximum iterations to 20, Lagrange solver precision tolerance to 2e-4, and remove weight decay for all hidden 2D weights, as retraction maintains the weight constraint (Section˜3.3).
-Training Data
+Training Data.
 We use the OLMo-Mix-1124 dataset (olmo20252olmo2furious), tokenized with the OLMo-2 tokenizer.
@@
 The router is implemented in FP32 precision.
-We adopt the auxiliary-loss-free strategy (wang2024auxiliarylossfreeloadbalancingstrategy), which demonstrated superior expert loading balance compared to global auxiliary loss (qiu2025demonsdetailimplementingload) in our ablations.
+We adopt the auxiliary-loss-free strategy (wang2024auxiliarylossfreeloadbalancingstrategy), which demonstrated superior expert load balancing compared to global auxiliary loss (qiu2025demonsdetailimplementingload) in our ablations.
 We enable expert bias with an update rate of 0.001.
-The sequence-level auxiliary loss is removed, as we found it redundant when expert bias is enabled.
+The sequence-level auxiliary loss is removed, as we find it redundant when expert bias is enabled.
 we use a sigmoid gate with top-$k$ scores renormalized and scaled by 2.
@@
 We sincerely thank Jianlin Su for his excellent blog, Scientific Spaces.
+We also thank Yifei Shen for his thoughtful discussions, and Franz Louis Cesista for helping clarify several conceptual formulations.
 Appendix A Appendix
@@
 We scale the hidden size, intermediate size, and number of attention heads (while fixing head dimensions).
-The models are trained on 30B tokens sampled from the Olmo2-mix124 dataset [olmo20252olmo2furious].
+The models are trained on 30B tokens sampled from the Olmo2-mix-1124 dataset [olmo20252olmo2furious].
 All optimizers use the Spectral $\bm{\mu}$P LR Scaler.